Light that is nonpolarized has an intensity of 6.0 W/m^2. Rays of this light are incident on two polarizing elements whose transmission axes are at an angle of 45°. What is the intensity of light emerging from the second film?



a. Given:


b. Formula:


c. Solution:


d. Final answer with unit:

1
Expert's answer
2022-05-04T17:38:30-0400

a.

"I_0=6~\frac W{m^2},\\
\theta=45°."

b.

"I=I_0\cos^2\theta."

c.

"I=6\cdot(\frac{\sqrt 2 }{2})^2=3."

d.

"3~\frac W{m^2}."